OpenStudy (anonymous):

:/ Division using synthetic division: g(x)=x^3-x^2-4x-6 What is the remainder?

OpenStudy (campbell_st):

what is your divisor.... and you can use the remainder theorem to find the remainder

OpenStudy (anonymous):

The problem states: Use synthetic division to determine whether or not x + 2 (a linear function) is a factor of (shown above) . The remainder in the division is?

OpenStudy (campbell_st):

it doesn't need synthetic division as g(-2) will tell you if x + 2 is a factor ... else it will give the remainder g(-2) = (-2)^3 -(-2)^2 -4(-2) - 6 = -8 -4 +8 - 6 g(-2) = -10 x+2 is not a factor and gives a remainder of -10 good luck with the division

OpenStudy (anonymous):

Oh I see. Thanks so much cambell_st.

OpenStudy (campbell_st):

oops the remainder is -10/(x+2)